OpenStudy (anonymous):

Jan has two jobs. She earns $8.90 per hour as a office clerk and $7.50 per hour as a teaching assistant. In one week, she works 12 hours at her office job and 15 hours at her teaching job. How much does Jan earn in one week?

OpenStudy (anonymous):

$219.30

OpenStudy (anonymous):

Those are some pretty bad jobs.

OpenStudy (anonymous):

7.50*15=112.50 8.90*12=106.80 106.80+112.50=219.30

OpenStudy (anonymous):

Are 5 people really looking at an answered question?